If Joe were 2 years older, he would be twice as old as Gary. Nicole at six years of age is half as old as Gary. How old is Joe?
Answer by addingup(3677)

J+2 = 2G
Nicole at 6 is 1/2 G, so 2(6) = G; 12 = G Substitute above:
J+2 = 2(12)
J+2 = 24
J = 22 Joe is 22
Nicole is 6
Gary is 12
Joe is 22